# PosIR: Position-Aware Heterogeneous Information Retrieval Benchmark

### v1

@@ Abstract
While dense retrieval models have achieved remarkable success, rigorous evaluation of their sensitivity to the position of relevant information (i.e., position bias) remains largely unexplored. Existing benchmarks typically employ position-agnostic relevance labels, conflating the challenge of processing long contexts with the bias against specific evidence locations. To address this challenge, we introduce PosIR (Position-Aware Information Retrieval), a comprehensive benchmark designed to diagnose position bias in diverse retrieval scenarios. PosIR comprises 310 datasets spanning 10 languages and 31 domains, constructed through a rigorous pipeline that ties relevance to precise reference spans, enabling the strict disentanglement of document length from information position. Extensive experiments with 10 state-of-the-art embedding models reveal that: (1) Performance on PosIR in long-context settings correlates poorly with the MMTEB benchmark, exposing limitations in current short-text benchmarks; (2) Position bias is pervasive and intensifies with document length, with most models exhibiting primacy bias while certain models show unexpected recency bias; (3) Gradient-based saliency analysis further uncovers the distinct internal attention mechanisms driving these positional preferences. In summary, PosIR serves as a valuable diagnostic framework to foster the development of position-robust retrieval systems. [^1] https://github.com/Ziyang1060/PosIR
PosIR: Position-Aware Heterogeneous Information Retrieval Benchmark
Ziyang Zeng1   Dun Zhang2   Yu Yan1 Xu Sun3Yudong Zhou2Yuqing Yang1thanks:   Corresponding author. 1Beijing University of Posts and Telecommunications 2Prior Shape 3Université Caen Normandie, ENSICAEN, CNRS, Normandie Univ, GREYC UMR6072, F-14000 Caen, France ziyang1060@bupt.edu.cn, dunnzhang0@gmail.com, yanyu2023@bupt.edu.cn xu.sun@unicaen.fr, zhouyudong@priorshape.com, yangyuqing@bupt.edu.cn
@@ 1 Introduction
The rapid advancement of dense retrieval has largely outpaced the evolution of rigorous evaluation protocols. While modern embedding models demonstrate impressive capabilities, characterizing their fine-grained behaviors requires benchmarks designed with higher diagnostic precision. In recent years, the Information Retrieval (IR) community has benefited from a series of milestone benchmarks. Early contributions like MS MARCO Nguyen et al. (2016) establish baselines for monolingual passage retrieval. Recognizing the linguistic diversity of the web, datasets such as MIRACL Zhang et al. (2023) extend the evaluation to multilingual scenarios. Additionally, benchmarks like BEIR Thakur et al. (2021), MMTEB Enevoldsen et al. (2025) and AIR-Bench Chen et al. (2025) aggregat diverse tasks to test general-purpose embeddings. These efforts have been instrumental in advancing the field.
@@
However, a critical dimension remains largely unexplored in existing benchmarks: position bias in information retrieval Hofstätter et al. (2021a). Specifically, current benchmarks generally operate on a position-agnostic assumption, providing relevance labels without specifying the location of the evidence. This conflation creates a diagnostic blind spot: when a retrieval model fails on a document, it is unclear whether the failure stems from the sheer volume of context (capacity limitation) or from bias against the location of the relevant span (positional sensitivity). Although Coelho et al. (2024); Zeng et al. (2025b) have observed primacy bias [^2] Primacy bias refers to models disproportionately focusing on the beginning of documents while overlooking relevant content appearing later. in embedding-based retrieval models, their experimental setups lack sufficient granularity to disentangle document length from information position. Without a controlled evaluation that orthogonalizes these variables, position bias cannot be strictly isolated from the model’s capacity to handle long contexts. Furthermore, the extent to which such bias varies across multilingual and cross-lingual retrieval settings remains an open question.
To address these gaps, we introduce PosIR (Position-Aware Information Retrieval), a comprehensive benchmark designed to rigorously evaluate and analyze position bias, which is characterized by three distinctive features: 1) Position-Aware: PosIR goes beyond coarse-grained document labels by associating each query with a precise reference span. By verifying relevance through strict span-based contrast, we enable the first systematic quantitative analysis of how the physical location of information impacts retrieval performance. 2) Heterogeneous: PosIR is designed to test robustness across diverse linguistic and semantic landscapes. It covers 31 domains and 10 languages (English, Chinese, and 8 translated languages), resulting in 310 datasets. This heterogeneity ensures that our findings on position bias are universal and not artifacts of a specific language or topic. 3) Length-Diverse: PosIR adopts a length-aware sampling strategy that spans documents from short passages to long texts. [^3] Up to 2048 tokens for native English and Chinese, with longer sequences for translated languages. This design enables the disentanglement of length and position factors, and exposes critical model behaviors, such as performance degradation in long-context settings, which are often overlooked by short-text benchmarks.
Using PosIR, we conduct extensive experiments on 10 popular IR models, yielding valuable insights into their performance characteristics:
  • Benchmarking Discrepancy: We find that retrieval performance on MMTEB correlates poorly with long-context retrieval performance on PosIR. Models that excel in short-context retrieval tasks often degrade significantly as document length increases.
  • Prevalence of Bias: Our evaluation reveals that position bias is pervasive and becomes more pronounced as document length increases, across both multilingual and cross-lingual retrieval settings. While most models exhibit primacy bias, favoring information at the beginning of documents, we identify unexpected recency bias in certain models.
  • Mechanistic Origins: Going beyond performance metrics, we employ gradient-based saliency analysis to uncover the internal mechanisms driving these biases. We identify two distinct attention behaviors that directly correlate with embedding models’ positional preferences over input tokens.
@@ 2 Benchmark Construction
The entire data generation pipeline of PosIR consists of four stages: 1) Bilingual corpora preparation, 2) Position-aware candidate generation, 3) Quality control, and 4) Multilingual translation. For detailed information of the models appearing in this paper, please refer to Table 12.
@@ 2.1 Bilingual Corpora Preparation
[…]
@@ 2.2 Position-Aware Candidate Generation
A retrieval dataset typically consists of three components: a document corpus, a set of queries, and relevance judgments (i.e., qrels). [^7] In TREC parlance, qrels refer to relevance judgments that indicate whether a document is relevant to a given query. After preparing the corpus in the initial stage, the candidate generation stage constructs the remaining two components of the dataset—queries and qrels—while explicitly accounting for the positional nature of relevance. As shown in Figure 1, the candidate generation process for a given domain proceeds as follows: 1) Sample one document from the corpus $\mathcal{D}$ as the positive document $d^{+}$. 2) Prompt LLMs to generate multiple queries $\{q_{i}\}_{i=1}^{m}$ for $d^{+}$ under a randomly sampled positional constraint (see Appendix C.1). 3) Prompt LLMs to locate the reference span $\mathrm{ref}_{i}$ corresponding to each generated question $q_{i}$ (see Appendix C.2). 4) Discard queries for which reference span localization fails (e.g., no unambiguous span can be identified or multiple candidate spans exist), and apply regular-expression matching to determine the character-level position span $\mathrm{pos}_{i}(\mathrm{start,end})$ of $\mathrm{ref}_{i}$ within $d^{+}$. In this stage, we use DeepSeek-V3.1 (Thinking Mode) DeepSeek-AI et al. (2025) for Steps 2 and 3, as these steps require multi-step reasoning over document content. After repeating the above steps thousands of times, we obtain the query set $\mathcal{Q}=\{q_{j}\}_{j=1}^{|\mathcal{Q}|}$, the positive document set $\mathcal{D}_{+}=\{d_{k}^{+}\}_{k=1}^{|\mathcal{D}_{+}|}$, the position-aware fine-grained qrels $\mathcal{R}_{+}=\{q_{j},d_{j}^{+},\mathrm{ref}_{j},\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})\}_{j=1}^{|\mathcal{Q}|}$.
@@ 2.3 Quality Control
Since all reference spans in the qrels $\mathcal{R}_{+}$ are generated by LLMs, they may suffer from localization errors or hallucinations. To identify and remove such cases, we evaluate the necessity of each reference span using relevance contrast. Specifically, we employ two re-ranking models (e.g., bge-reranker-v2-minicpm-layerwise and Qwen3-Reranker-4B Zhang et al. (2025)) to compute two relevance scores for each query $q_{j}$: $s^{w/}_{j}$, the relevance score between $q_{j}$ and the full positive document $d_{j}^{+}$, and $s^{w/o}_{j}$, the relevance score between $q_{j}$ and a modified document obtained by removing the reference span $\mathrm{ref}_{j}$ from $d_{j}^{+}$. We retain only those queries for which $s^{w/}_{j}>s^{w/o}_{j}$ under both re-ranking models, indicating that the reference span is indispensable for establishing query-document relevance. Furthermore, for the remaining queries, we perform an additional verification step using an LLM-based relevance assessor (see Appendix C.3). Following the same contrastive setup, the LLM produces two 5-level relevance labels, $s^{\prime w/}_{j}$ and $s^{\prime w/o}_{j}$, ranging from 0 (completely irrelevant) to 4 (perfectly relevant). We retain a query only if $s^{\prime w/}_{j}>s^{\prime w/o}_{j}$ and $s^{\prime w/}_{j}\geq 3$, corresponding to a highly relevant judgment. Through this filtering process, we ensure that each query-document pair is relevant specifically due to a well-defined reference span at a particular position. This property is essential for accurately studying position bias in retrieval models.
@@
After applying the above quality control process to each query, we obtain the refined query set $\mathcal{Q}^{\prime}$, corpus $\mathcal{D}^{\prime}$, and qrels $\mathcal{R}_{+}^{\prime}$, which together constitute the final bilingual dataset for each domain. Overall, this intentionally conservative quality control process prioritizes annotation precision over coverage, which is essential for isolating and analyzing fine-grained positional effects.
@@ 2.4 Multilingual Translation
After completing the earlier stages, we obtain high-quality bilingual English-Chinese datasets covering 31 domains, resulting in 62 position-aware retrieval datasets. To comprehensively study position bias in both multilingual and cross-lingual retrieval tasks, it is essential to consider a broad range of language scenarios. LLMs have exhibited remarkable proficiency in multilingual machine translation Zhu et al. (2024), significantly improving efficiency and eliminating the need to manually create position-aware retrieval datasets for each language from scratch. We employ Qwen3-30B-A3B-Instruct-2507 Yang et al. (2025a) to translate the English retrieval datasets across 31 domains into 8 languages, such as French, Spanish, Russian, and others, resulting in 248 datasets. Specifically, for each domain, we translate the English datasets $\mathcal{Q}_{en}^{\prime}$ and $\mathcal{D}_{en}^{\prime}$ into the target language datasets $\mathcal{Q}_{*}^{\prime}$ and $\mathcal{D}_{*}^{\prime}$, while the qrels $\mathcal{R}_{+}^{\prime}$ are shared across languages. This means that the position-aware relevance labels in the other languages remain consistent with the English dataset. To evaluate the translation quality, we employ both LLM-based automatic evaluation Kocmi and Federmann (2023) and human evaluation on a sampled subset, both of which demonstrate the overall high quality of the translations. However, during the multilingual translation process, we observe that token lengths grow disproportionately across languages due to richer morphology and longer sentence structures, which can cause the translated retrieval corpus $\mathcal{D_{*}}^{\prime}$ to exceed the predefined 2048-token length limit. Additionally, we observe issues such as invalid prefixes, refusal to translate, and repetitive or verbose translations. To further improve the translation quality, we implement a set of fixing and filtering strategies to ensure clarity and consistency in the results. For more details on the multilingual machine translation stage, please refer to Appendix B.

### v2

@@ Abstract
In real-world documents, the information relevant to a user query may reside anywhere from the beginning to the end. This makes position bias—a systematic tendency of retrieval models to favor or neglect content based on its location—a critical concern. Although recent studies have identified such bias, existing analyses focus predominantly on English, fail to disentangle document length from information position, and lack a standardized framework for systematic diagnosis. To address these limitations, we introduce PosIR (Position-Aware Information Retrieval), the first standardized benchmark designed to systematically diagnose position bias in diverse retrieval scenarios. PosIR comprises 310 datasets spanning 10 languages and 31 domains, with relevance tied to precise reference spans. At its methodological core, PosIR employs a length-controlled bucketing strategy that groups queries by positive document length and analyzes positional effects within each bucket. This design strictly isolates position bias from length-induced performance degradation. Extensive experiments on 10 state-of-the-art embedding-based retrieval models reveal that: (1) retrieval performance on PosIR with documents exceeding 1536 tokens correlates poorly with the MMTEB benchmark, exposing limitations of current short-text evaluations; (2) position bias is pervasive in embedding models and even increases with document length, with most models exhibiting primacy bias while certain models show unexpected recency bias; (3) as an exploratory investigation, gradient-based saliency analysis further uncovers two distinct internal mechanisms that correlate with these positional preferences. We hope that PosIR can serve as a valuable diagnostic framework to advance the development of position-robust retrieval systems. [^1] https://github.com/Ziyang1060/PosIR
@@ 1 Introduction
Information retrieval aims to identify relevant documents from a large collection in response to a user query [17]. A robust retrieval system should assess a document’s relevance based on its content, not on where that content happens to appear. However, retrieval models that systematically favor or neglect content based on its location risk missing critical information and undermining retrieval quality. This phenomenon, known as position bias [10], has attracted growing attention: recent studies [5, 9, 33] have empirically shown that embedding-based retrieval models exhibit primacy bias, disproportionately attending to the early portions of documents while neglecting evidence that appears later.
Although position bias has been recognized, existing analyses suffer from three critical limitations. First, the linguistic scope is narrow. Prior investigations focus almost exclusively on English, leaving open the question of whether position bias generalizes across typologically diverse languages and how it manifests in cross-lingual retrieval settings. Second, document length and information position are confounded. Current studies typically vary the position of relevant information without controlling for document length [33]. As a result, when a model performs poorly on a document with late-appearing evidence, it remains ambiguous whether the failure reflects a genuine positional preference or simply a capacity limitation when processing long documents. Without an evaluation design that isolates these two variables, position bias cannot be rigorously measured. Third, a standardized evaluation framework is absent. Different studies adopt varying document lengths, position definitions, and analytical protocols, making their conclusions difficult to compare and reproduce. The lack of a unified benchmark prevents the community from systematically tracking progress or diagnosing model-specific weaknesses.
To address these limitations, we introduce PosIR (Position-Aware Information Retrieval), the first standardized benchmark for systematically diagnosing position bias in information retrieval. PosIR is characterized by three key design principles, which collectively address the limitations identified above: 1) Heterogeneous Coverage: PosIR spans 10 languages (English, Chinese, and 8 translated languages) and 31 domains, yielding 310 datasets in total. Both automatic and human quality assessments confirm that the translated data faithfully preserves source-language semantics, ensuring that our findings generalize beyond English and are not artifacts of a specific language. 2) Length-Controlled Analysis: At its methodological core, PosIR employs a length-controlled bucketing strategy: queries are grouped by the positive document length, and positional effects are analyzed within each bucket. This design strictly isolates position bias from length-induced performance degradation, directly addressing the length–position confound. 3) Position-Aware Relevance: Unlike existing benchmarks that assign coarse document-level relevance labels [18, 26, 4], PosIR associates each query with a precise reference span within the document. Relevance is verified through strict span-based contrast, enabling fine-grained quantitative analysis of how the physical location of information impacts retrieval performance.
Using PosIR, we conduct extensive experiments on 10 popular embedding-based retrieval models, yielding several key insights:
  • Benchmarking Discrepancy. Retrieval performance on PosIR with documents exceeding 1,536 tokens correlates poorly with the MMTEB [8] scores. Models that excel in short-context evaluations often degrade significantly as document length increases.
  • Prevalence of Bias. Position bias is pervasive in embedding models and even increases with document length, across both multilingual and cross-lingual settings. While most models exhibit primacy bias, we first identify unexpected recency bias in NV-Embed-v2 [15].
  • Mechanistic Origins. Going beyond performance metrics, we employ an exploratory gradient-based saliency analysis to uncover the internal mechanisms driving these biases, identifying two distinct attention behaviors that correlate with embedding models’ positional preferences over input tokens.
2 Related Work
2.1 Position Bias in Information Retrieval
Prior work has widely documented the existence of position bias in neural information retrieval (IR) systems. Previous work [11] identified that MS MARCO [19], a foundational IR dataset, is heavily skewed toward head-located relevant spans, leading fine-tuned models to inherit and amplify this position bias [12], compromising the reliability of existing evaluations [23, 25]. Recent studies confirm that such bias persists even in state-of-the-art embedding models [5, 9, 33]. Specifically, previous work [33] utilized answer start positions in SQuAD [22] for bias analysis; however, SQuAD is constrained to English monolingual data and brief passages (avg. 117 words), hindering the generalization of findings to multilingual or long-context scenarios. Despite these insights, a standardized benchmark to rigorously evaluate position bias, especially within multilingual and longer contexts, remains absent. Consequently, PosIR is proposed to bridge this gap.
2.2 Synthetic Data for Information Retrieval
Leveraging large language models (LLMs) for synthetic data generation has recently attracted growing attention in the IR community. Existing works can be broadly categorized into two primary use cases: training and evaluation. From a training perspective, LLMs are employed to address the scarcity of domain-specific or task-specific data [34]. Methods like InPars [2] and Promptagator [6] prompt LLMs to generate synthetic queries for specific documents, facilitating effective retriever training in zero-shot settings. This paradigm has been further extended to support multilingual retrieval tasks [27, 29]. Regarding evaluation, recent studies have validated the reliability of synthetic test collections for benchmarking [21]. LLM-generated pseudo-labels and queries have been leveraged for unsupervised model selection on a target corpus [13]. Additionally, frameworks like AIR-Bench [4] provide automated pipelines to generate evaluation data for emerging domains efficiently. Inspired by these advancements, we extend the application of synthetic data generation to the specific challenge of position bias. Unlike prior works that focus on document-level relevance, we utilize LLMs to generate fine-grained, position-aware relevance annotations, which are crucial for analyzing position bias. By incorporating rigorous quality control mechanisms, PosIR ensures the reliability of synthetic signals, enabling efficient and effective diagnosis of position bias in retrieval models.
@@ 3 Benchmark Construction
The entire data generation pipeline of PosIR consists of four stages: 1) Bilingual corpora preparation, 2) Position-aware candidate generation, 3) Quality control, and 4) Multilingual translation.
@@ 3.1 Bilingual Corpora Preparation
As illustrated in Figure 1, the first stage involves preparing high-quality bilingual corpora across diverse domains. Specifically, we leverage IndustryCorpus2222https://huggingface.co/datasets/BAAI/IndustryCorpus2, an open-source and carefully curated English-Chinese dataset that covers 31 industry categories, broadly representing mainstream real-world application domains. We exclude the original Others category to maintain domain specificity and introduce FineWeb to represent the general domain: for English, we use the fineweb sampled 10BT subset [^3] https://huggingface.co/datasets/HuggingFaceFW/fineweb/tree/main/sample/10BT, while for Chinese we adopt Fineweb-Edu-Chinese-V2.1444https://huggingface.co/datasets/opencsg/Fineweb-Edu-Chinese-V2.1. After this adjustment, the resulting corpus still consists of 31 domains. Then, we adopt a length-aware sampling strategy that moderately over-samples shorter documents, which are more prevalent in real-world corpora. Specifically, using the Qwen3 tokenizer [30], we partition documents into eight length buckets with an interval of 256 tokens (covering 0 to 2048 tokens) and apply a sampling ratio of 3:3:3:3:2:2:2:2, such that documents under 1024 tokens are sampled at 1.5 $\times$ the rate of longer ones. Following this strategy, we sample approximately 70k documents for each domain from both the English and Chinese corpora. Formally, for a given domain, we denote the English and Chinese corpora as $\mathcal{D}_{en}=\{d_{i}\}_{i=1}^{n_{en}}$ and $\mathcal{D}_{zh}=\{d_{i}\}_{i=1}^{n_{zh}}$, containing $n_{en}$ and $n_{zh}$ documents, respectively. More generally, we use $\mathcal{D}=\{d_{i}\}_{i=1}^{n}$ to denote a corpus in an arbitrary language, where $n$ denotes the number of documents in that corpus.
@@ 3.2 Position-Aware Candidate Generation
A retrieval dataset typically consists of three components: a document corpus, a set of queries, and relevance judgments (i.e., qrels). [^5] In TREC parlance, qrels refer to relevance judgments that indicate whether a document is relevant to a given query. After preparing the corpus in the initial stage, the candidate generation stage constructs the remaining two components of the dataset—queries and qrels—while explicitly accounting for the positional nature of relevance. As shown in Figure 1, the candidate generation process for a given domain proceeds as follows: 1) Sample one document from the corpus $\mathcal{D}$ as the positive document $d^{+}$. 2) Prompt LLMs to generate multiple queries $\{q_{i}\}_{i=1}^{m}$ for $d^{+}$ under a randomly sampled positional constraint (see Appendix 0.D.1). 3) Prompt LLMs to locate the reference span $\mathrm{ref}_{i}$ corresponding to each generated question $q_{i}$ (see Appendix 0.D.2). 4) Discard queries for which reference span localization fails (e.g., no unambiguous span can be identified, or multiple candidate spans exist), and apply regular-expression matching to determine the character-level position span $\mathrm{pos}_{i}(\mathrm{start,end})$ of $\mathrm{ref}_{i}$ within $d^{+}$. In this stage, we use DeepSeek-V3.1 (Thinking Mode) [7] for Steps 2 and 3, as these steps require multi-step reasoning over document content. After repeating the above steps thousands of times, we obtain the query set $\mathcal{Q}=\{q_{j}\}_{j=1}^{|\mathcal{Q}|}$, the positive document set $\mathcal{D}_{+}=\{d_{k}^{+}\}_{k=1}^{|\mathcal{D}_{+}|}$, the position-aware fine-grained qrels $\mathcal{R}_{+}=\{q_{j},d_{j}^{+},\mathrm{ref}_{j},\mathrm{pos}_{j}(\mathrm{start},\mathrm{end})\}_{j=1}^{|\mathcal{Q}|}$.
@@ 3.3 Quality Control
Since all reference spans in the qrels $\mathcal{R}_{+}$ are generated by LLMs, they may suffer from localization errors or hallucinations. To identify and remove such cases, we evaluate the necessity of each reference span using relevance contrast. Specifically, we employ two re-ranking models (e.g., bge-reranker-v2-minicpm-layerwise and Qwen3-Reranker-4B [37]) to compute two relevance scores for each query $q_{j}$: $s^{w/}_{j}$, the relevance score between $q_{j}$ and the full positive document $d_{j}^{+}$, and $s^{w/o}_{j}$, the relevance score between $q_{j}$ and a modified document obtained by removing the reference span $\mathrm{ref}_{j}$ from $d_{j}^{+}$. We retain only those queries for which $s^{w/}_{j}>s^{w/o}_{j}$ under both re-ranking models, indicating that the reference span is indispensable for establishing query-document relevance. For the queries that pass this initial filter, we perform a second verification step using an LLM-based relevance assessor (see Appendix 0.D.3). Following the same contrastive setup, the LLM produces two 5-level relevance labels, $s^{\prime w/}_{j}$ and $s^{\prime w/o}_{j}$, ranging from 0 (completely irrelevant) to 4 (perfectly relevant). We retain a query only if $s^{\prime w/}_{j}>s^{\prime w/o}_{j}$ and $s^{\prime w/}_{j}\geq 3$, corresponding to a highly relevant judgment. We acknowledge that removing the reference span may disrupt local coherence, potentially causing relevance scores ($s^{w/o}$) to decrease even when the span is not truly indispensable. To mitigate this, the LLM-based verification assesses relevance at a semantic level, providing a complementary check that is less susceptible to surface-level fluency disruption. Through this dual verification process, we ensure that each retained query-document pair is relevant specifically due to a well-defined reference span at a particular position—a property essential for accurately studying position bias.
@@
After applying the above quality control process to each query, we obtain the refined query set $\mathcal{Q}^{\prime}$, corpus $\mathcal{D}^{\prime}$, and qrels $\mathcal{R}_{+}^{\prime}$, which together constitute the final bilingual dataset for each domain. Overall, this intentionally conservative quality control process prioritizes annotation precision over coverage, which is essential for isolating and analyzing fine-grained positional effects. To further validate data quality, we reviewed 500 generated English query–document–reference span triples (details in Appendix 0.B), confirming that our automated pipeline introduces minimal risk of accumulated hallucinations or semantic drift.
@@ 3.4 Multilingual Translation
After completing the earlier stages, we obtain 62 high-quality English and Chinese datasets covering 31 domains. To extend linguistic coverage, we employ Qwen3-30B-A3B-Instruct-2507 [30] to translate the English datasets across 31 domains into 8 additional languages (French, Spanish, Russian, etc.), yielding 248 datasets. Specifically, for each domain, we translate the English datasets $\mathcal{Q}_{en}^{\prime}$ and $\mathcal{D}_{en}^{\prime}$ into the target language datasets $\mathcal{Q}_{*}^{\prime}$ and $\mathcal{D}_{*}^{\prime}$, while the qrels $\mathcal{R}_{+}^{\prime}$ are shared across source and translated languages to ensure consistent position-aware relevance labels. While character-level span indices inevitably shift across languages due to differences in morphology and syntax, our position bias analysis operates on coarse-grained relative positions, making it robust to such minor deviations. To evaluate translation quality, we employ both LLM-based automatic evaluation [14] and human evaluation on a sampled subset, both confirming overall high quality. However, we observe that token lengths grow disproportionately across languages, sometimes causing the translated corpus to exceed the predefined 2048-token limit. We also encounter issues such as invalid prefixes, refusal to translate, and verbose outputs. To address these issues and further improve translation quality, we implement a set of fixing and filtering strategies for translated languages. For more details on the multilingual machine translation stage, please refer to Appendix 0.C.